OpenStudy (anonymous):

PLEASE HELP. VERY DESPERATE suppose you deposit $1,000 in a savings account that pays interest at an annual rate no money is added or withdrawn from the account.

OpenStudy (anonymous):

Where is the question?

OpenStudy (anonymous):

what will the amount be after 4 years

OpenStudy (anonymous):

You can use the formula for a simple interest rate which is interest = principal * rate * time You will need to know the interest rate though.

OpenStudy (anonymous):

the interest rate is 6%

OpenStudy (anonymous):

I = 1,000 * 0.06 * 4 = 240 so there will be $1,240 after 4 years.

OpenStudy (anonymous):

what about after 17 years?

OpenStudy (anonymous):

Just substitute 17 in for 4.

OpenStudy (anonymous):

i dont get it? :(

OpenStudy (anonymous):

You have the initial amount which is called the principal. You then multiply that by the amount of interest earned, 6%, converted into decimal form which is 0.06. Then multiply that by the amount of time.

OpenStudy (anonymous):

For 17 years it would be I = 1000 * 0.06 * 17 = 1020 Add that to the principal and you have your total after 17 years. 1020+ 1000 = 2020

OpenStudy (anonymous):

for the account to contain 1,500 how long would it take?

OpenStudy (anonymous):

Solve this equation: 1000 * 0.06 * x = 1,500

OpenStudy (anonymous):

type a whole number

OpenStudy (anonymous):

8 years

OpenStudy (anonymous):

what about 2,000

OpenStudy (anonymous):

16 years 8 months

OpenStudy (anonymous):

kk thanks!:)

OpenStudy (anonymous):

ps. YOUR A flutterING IDIOT WHO MADE ME GET IT WRONG

OpenStudy (anonymous):

READ A MATH BOOK IDIOT